# On equationally Noetherian predicate structures

### v1

@@ 1. Introduction
(1) $\mathcal{A}\not\models s_{i}(a_{i})\text{ for all }i\in\mathbb{N}\text{, and }\mathcal{A}\models s_{j}(a_{i})\text{ for all }j<i.$
@@ 3.1. Configurations of predicate equations
Proof.
Let $S=S_{C}\cup S_{X}\cup S_{X,C}$, where the subsystem $S_{X,C}$ consists of equations that have variables and constants at the same time, $S_{X}$ consists of equations with no constants, and $S_{C}$ consists of equations with no variables. Because the number of variables is finite, the subsystem $S^{X}$ is finite. The fact that $S_{C}$ is finite follows from the assumption of the proposition. Therefore, the subsystem $S_{X,C}$ is infinite. Let $S_{X,C}=S_{P_{1}}\cup S_{P_{2}}\cup\ldots\cup S_{P_{m}}\cup S_{=}$, where $S_{P_{i}}$ consists of equations of the form $P_{i}(w_{1},\ldots,w_{n_{i}})$, and $S_{=}$ consists of $w_{1}=w_{2}$. Since $S_{X,C}$ is infinite, at least one of the subsystems $S_{P_{1}},S_{P_{2}},\ldots,S_{P_{m}},S_{=}$ is infinite. ∎
@@
Proof.
Since $\mathcal{A}$ is not equationally Noetherian, it follows from Corollary 1 that there are sequences of elements $(a_{i})_{i\in\mathbb{N}}$ and of equations $S(X)$ such that (1) holds. From Proposition 1, there is an infinite subsystem $S^{\prime}\subseteq S$ such that all its equations contain one predicate $P^{(n)}$ of $\mathcal{L}_{\mathcal{A}}$, i.e., $S^{\prime}=\{P(w_{1}^{i},\ldots,w_{n}^{i})\}_{i\in\mathbb{N}}$, where $w_{j}^{i}\in X\cup A$. According to Remark 2, there is an infinite subsystem $S^{\prime\prime}\subseteq S^{\prime}$ such that it is not equivalent to any of its finite subsystems and consists of predicate equations of the same configuration. ∎
@@ 3.2. Projections of predicates and gluings of predicates
Example 2.
Consider the hypergraph $\Gamma$ from Example 1. Let $I=\{\{1\},\{2,3\}\}$ be an exact partition of the set $\{1,2,3\}$. Then $\nicefrac{{E}}{{I}}$ is the binary predicate that is true only for pairs $(v_{1},v_{3}),(v_{3},v_{2})$.
@@ 3.3. Perfectly non-Noetherian structures
This notion is similar to the notion of a perfectly non-Noetherian substructure introduced above. The difference is that the sequences of elements $\{a_{i}\}_{i\in\mathbb{N}}$ and $\{b_{i}\}_{i\in\mathbb{N}}$ from Definition 3 coincide.
@@ 3.4. Criterion for a predicate structure to be equationally Noetherian
If $n>2$, then we perform the procedure described above for all the pairs of columns. At each step for a selected pair of columns in the current subsequence $D^{\prime}$ of $D_{n}$, we choose a new subsequence $D^{\prime\prime}$ and, for the next pair of columns, we perform the procedure on the subsequence $D^{\prime\prime}$. After performing the procedure on all the pairs of columns, we obtain a subsequence $\tilde{D}$ that satisfies the condition of the lemma. ∎
@@
Let the arity of the predicate $P_{\tilde{D},I}$ be greater than $1$. Then, all the elements from $\tilde{D}$ define a $P_{\tilde{D},I}$perfectly non-Noetherian substructure, and, therefore, by Definition 3, the structure $\mathcal{A}$ has a $P_{\tilde{D},I}$-perfectly non-Noetherian substructure.
@@
The case when an algebraic structure contains a non-Noetherian clique follows from Remark 3. ∎
@@
Proof.
Since $\mathcal{A}$ is not equationally Noetherian, it follows from Corollary 1 that there are sequences of equations $S(X)$ and elements $\{a_{i}\}_{i\in\mathbb{N}}$ satisfying Lemma 1. From Proposition 1, there is an infinite subsystem of equations $S_{P}\subseteq S$ written with only one predicate symbol $P$ of the language $\mathcal{L}_{\mathcal{A}}$ such that it is not equivalent to any of its subsystems because of Corollary 1. Let $\mathcal{L}_{\mathcal{A}}^{\prime}$ consist of $P$ and all the constant symbols from $A$. Note that the system $S_{P}$ can be considered over the algebraic structure $\mathcal{A}^{\prime}=\langle A,\mathcal{L}_{\mathcal{A}}^{\prime}\rangle$, where the interpretations of the symbols of the language $\mathcal{L}_{\mathcal{A}}^{\prime}$ coincide with the corresponding interpretations of these symbols in the original algebraic structure $\mathcal{A}$. Then $\mathcal{A}^{\prime}$ is not equationally Noetherian. ∎
Lemmas 4 and 5, in the form of the following theorem, are a criterion for arbitrary predicate structures to be equationally Noetherian in terms of forbidden substructures.
@@ 4. Equationally Noetherian graphs, hypergraphs, and partial orders
Proof.
The equivalence $1\iff 2$ follows from Theorem 1. The equivalence $1\iff 3$ follows from Theorem 3. Therefore, $2\iff 3$. ∎
@@
If $B$ is a subset such that its upper base cone is infinitely generated, the reasoning is similar with the difference that it is necessary to use the predicate $\succ$. ∎
Example 4.
Consider the natural order of integers $\mathcal{Z}=\langle\mathbb{Z},\prec\rangle$. Note that the conditions from Definition 4 hold for the sequence $\{i\}_{i\in\mathbb{N}}$, i.e., the order $\mathcal{Z}$ contains a non-Noetherian clique. The sequences of elements $\{-2\cdot i\}_{i\in\mathbb{N}}$ and equations $\{x\prec-2\cdot i-1\}_{i\in\mathbb{N}}$, by Definition 3, define a perfectly non-Noetherian substructure of $\mathcal{Z}$.

### v3

@@ Abstract.
Key words and phrases:
predicate structures, equationally Noetherian structures, universal algebraic geometry
2020 Mathematics Subject Classification. 08A99, 03C05.
@@ 1. Introduction
$\mathcal{A}\not\models s_{i}(a_{i})\text{ for all }i\in\mathbb{N}\text{, and }\mathcal{A}\models s_{j}(a_{i})\text{ for all }j<i.$ (1)
@@ 3.1. Configurations of predicate equations
Proof 3.1.
Let $S=S_{C}\cup S_{X}\cup S_{X,C}$, where the subsystem $S_{X,C}$ consists of equations that have variables and constants at the same time, $S_{X}$ consists of equations with no constants, and $S_{C}$ consists of equations with no variables. Because the number of variables is finite, the subsystem $S^{X}$ is finite. The fact that $S_{C}$ is finite follows from the assumption of the proposition. Therefore, the subsystem $S_{X,C}$ is infinite. Let $S_{X,C}=S_{P_{1}}\cup S_{P_{2}}\cup\ldots\cup S_{P_{m}}\cup S_{=}$, where $S_{P_{i}}$ consists of equations of the form $P_{i}(w_{1},\ldots,w_{n_{i}})$, and $S_{=}$ consists of $w_{1}=w_{2}$. Since $S_{X,C}$ is infinite, at least one of the subsystems $S_{P_{1}},S_{P_{2}},\ldots,S_{P_{m}},S_{=}$ is infinite.
@@
Proof 3.2.
Since $\mathcal{A}$ is not equationally Noetherian, it follows from Corollary 2 that there are sequences of elements $(a_{i})_{i\in\mathbb{N}}$ and of equations $S(X)$ such that (1) holds. From Proposition 3, there is an infinite subsystem $S^{\prime}\subseteq S$ such that all its equations contain one predicate $P^{(n)}$ of $\mathcal{L}_{\mathcal{A}}$, i.e., $S^{\prime}=\{P(w_{1}^{i},\ldots,w_{n}^{i})\}_{i\in\mathbb{N}}$, where $w_{j}^{i}\in X\cup A$. According to Remark 5, there is an infinite subsystem $S^{\prime\prime}\subseteq S^{\prime}$ such that it is not equivalent to any of its finite subsystems and consists of predicate equations of the same configuration.
@@ 3.2. Projections of predicates and gluings of predicates
{exa}
Consider the hypergraph $\Gamma$ from Example 3.2. Let $I=\{\{1\},\{2,3\}\}$ be an exact partition of the set $\{1,2,3\}$. Then $\nicefrac{{E}}{{I}}$ is the binary predicate that is true only for pairs $(v_{1},v_{3}),(v_{3},v_{2})$.
@@ 3.3. Perfectly non-Noetherian structures
This notion is similar to the notion of a perfectly non-Noetherian substructure introduced above. The difference is that the sequences of elements $\{a_{i}\}_{i\in\mathbb{N}}$ and $\{b_{i}\}_{i\in\mathbb{N}}$ from Definition 3.3 coincide.
@@ 3.4. Criterion for a predicate structure to be equationally Noetherian
If $n>2$, then we perform the procedure described above for all the pairs of columns. At each step for a selected pair of columns in the current subsequence $D^{\prime}$ of $D_{n}$, we choose a new subsequence $D^{\prime\prime}$ and, for the next pair of columns, we perform the procedure on the subsequence $D^{\prime\prime}$. After performing the procedure on all the pairs of columns, we obtain a subsequence $\tilde{D}$ that satisfies the condition of the lemma.
@@
Let the arity of the predicate $P_{\tilde{D},I}$ be greater than $1$. Then, all the elements from $\tilde{D}$ define a $P_{\tilde{D},I}$perfectly non-Noetherian substructure, and, therefore, by Definition 3.3, the structure $\mathcal{A}$ has a $P_{\tilde{D},I}$-perfectly non-Noetherian substructure.
@@
The case when an algebraic structure contains a non-Noetherian clique follows from Remark 7.
@@
Proof 3.5.
Since $\mathcal{A}$ is not equationally Noetherian, it follows from Corollary 2 that there are sequences of equations $S(X)$ and elements $\{a_{i}\}_{i\in\mathbb{N}}$ satisfying Lemma 1. From Proposition 3, there is an infinite subsystem of equations $S_{P}\subseteq S$ written with only one predicate symbol $P$ of the language $\mathcal{L}_{\mathcal{A}}$ such that it is not equivalent to any of its subsystems because of Corollary 2. Let $\mathcal{L}_{\mathcal{A}}^{\prime}$ consist of $P$ and all the constant symbols from $A$. Note that the system $S_{P}$ can be considered over the algebraic structure $\mathcal{A}^{\prime}=\langle A,\mathcal{L}_{\mathcal{A}}^{\prime}\rangle$, where the interpretations of the symbols of the language $\mathcal{L}_{\mathcal{A}}^{\prime}$ coincide with the corresponding interpretations of these symbols in the original algebraic structure $\mathcal{A}$. Then $\mathcal{A}^{\prime}$ is not equationally Noetherian.
Lemmas 9 and 10, in the form of the following theorem, are a criterion for arbitrary predicate structures to be equationally Noetherian in terms of forbidden substructures.
@@ 4. Equationally Noetherian graphs, hypergraphs, and partial orders
Proof 4.1.
The equivalence $1\iff 2$ follows from Theorem 11. The equivalence $1\iff 3$ follows from Theorem 13. Therefore, $2\iff 3$.
@@
If $B$ is a subset such that its upper base cone is infinitely generated, the reasoning is similar with the difference that it is necessary to use the predicate $\succ$.
{exa}
Consider the natural order of integers $\mathcal{Z}=\langle\mathbb{Z},\prec\rangle$. Note that the conditions from Definition 3.3 hold for the sequence $\{i\}_{i\in\mathbb{N}}$, i.e., the order $\mathcal{Z}$ contains a non-Noetherian clique. The sequences of elements $\{-2i\}_{i\in\mathbb{N}}$ and equations $\{x\prec-2i-1\}_{i\in\mathbb{N}}$, by Definition 3.3, define a perfectly non-Noetherian substructure of $\mathcal{Z}$.